Tom has two loans totaling 18000.00 one loan has a interest of 4.25% the other loan is at 6% after six years tom has a total interest of 5400.00 what is the total of both loans
Answer by mananth(16946)

Loan A 4.25% per annum ----x
Loan B 6.00% per annum ----y
for 6 years interest interest = 5400
for 1 year it will be 900

x+y= 18000 ------------------------1
4.25%x+6.00%y=$900.00
Multiply by 100
4.25x+6y=$90,000.00 --------2

Multiply (1) by -4.25
we get

-4.25 x -4.25 y = -76500.00

Add this to (2)

0 x 1.75 y = $13,500.00

divide by 1.75

y = $7,714.29 investment at 6.00%
Balance $10,285.71 investment at 4.25%
CHECK
$10,285.71 $437.14
$7,714.29 $462.86
Total $900.00